OpenStudy (anonymous):

Graph the given equation. y = 5x − 5

OpenStudy (anonymous):

Put various values of x like -1, -2, 0, 1, 5 and find corresponding values of y from the given equation Plot the point and get the graph

OpenStudy (anonymous):

just plot a couple of point and write the line

OpenStudy (anonymous):

its a linear function in the book there is an example but i cant seem to follow it

OpenStudy (anonymous):

Of course it is a linear function

OpenStudy (anonymous):

You are trying to find point (X,Y) so for example if X = 0 (0,Y) what does Y=?

OpenStudy (anonymous):

Suppose you put x= 1 in the given eqn, you get y= 5(1) -5 = 5-5 = 0 So you get a point (1,0)

OpenStudy (anonymous):

then find many points and connect the dots :)

OpenStudy (anonymous):

\[y=5x-5\]

OpenStudy (anonymous):

ooh ok so i need to get many point and do a line right? [could you see the attached example?]

OpenStudy (anonymous):

the image file gives you the nature of the plot...

OpenStudy (anonymous):

Yes 3-4 point r more than enough then plot them and you will get the required graph

OpenStudy (anonymous):

oh great thnx! but i have one more question it the same kind of problem y-1/2x=-4 do i replace the x for different numbers and plug that in?

OpenStudy (anonymous):

or am i mistaken?

OpenStudy (anonymous):

Yes you can convert it to y = -4 + 1/2x Then put values for x and get values of y

OpenStudy (anonymous):

ok thnx alot!

OpenStudy (anonymous):

Medal for me ?? (:-)))

OpenStudy (anonymous):

how do i give one?

OpenStudy (anonymous):

Look on the right, You will see "Give Medal" icon in front of my name Just click on it

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

Strange, I can see that icon against your name at my end!!! By the way, how did you capture and send this screenshot?? Then I can send mine to show you

OpenStudy (anonymous):

For a linear equation you really only need two points to graph it.

OpenStudy (anonymous):

I am on a mac so i click command+shift+4 and select the section i want to give the screen shot. on pc i could click tue button at the top right that says PrintQ or something like that and crop it in pait.exe then send it

OpenStudy (anonymous):

crop it on paint.exe*

OpenStudy (anonymous):

Well, Im on a pc so cant do it, anyway it doesnt matter.. Your problem is resolved i hope :))

OpenStudy (anonymous):

yeah, for this one :X... I am gonna as another one on a way different topic, and sorry i couldnt give you a medal I will try to find out how to... (became a fan so if i figure out how to I will give it to you ;D)

OpenStudy (anonymous):

Sure, thx for becoming my fan.....